AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 213     Entry time: Sat Feb 19 20:08:43 2022
Author: JM, CB 
Subject: Beam ON run 49 

Run started 19/2/22 21:05. Beam ON

Temperatures OK

System wide checks OK - white rabbit fails on all.

Statistics are within normal range (10-20K goodevents)

Leakage current 3.6 uA

potentiometer = 20.28 mm

 

20/2/22

Leakage current ~3.675 uA. Looks stable now motors can not move (attachment 1)

potentiometer = 20.28 mm

Temperatures OK

System wide checks OK - white rabbit fails all

Statistic within normal range (see attachment 2 - ADC data items) Counter for aida 1 and aida 2 much higher due to noise bursts (elog 212)

 

XY total plot (attachment 3)

Low energy Ex histogram (attachment 4)

Low energy Ey histogram (attachment 5)

Tape Server data written (attachment 6)

 

Run end 20/2/22 15:08

Attachment 1: 2022-02-20_14-16-16-leakage-current.png  21 kB  Uploaded Sun Feb 20 13:16:44 2022  | Hide | Hide all
2022-02-20_14-16-16-leakage-current.png
Attachment 2: 2022-02-20_14-18-07-statistic-ADCdata.png  101 kB  Uploaded Sun Feb 20 13:18:36 2022  | Hide | Hide all
2022-02-20_14-18-07-statistic-ADCdata.png
Attachment 3: 2022-02-20_14-20-43-xyTotal.png  241 kB  Uploaded Sun Feb 20 13:21:07 2022  | Hide | Hide all
2022-02-20_14-20-43-xyTotal.png
Attachment 4: 2022-02-20_14-24-19-LowEnergy-ExTotal.png  196 kB  Uploaded Sun Feb 20 13:25:05 2022  | Hide | Hide all
2022-02-20_14-24-19-LowEnergy-ExTotal.png
Attachment 5: 2022-02-20_14-28-02-LowEnergy-EyTotal.png  196 kB  Uploaded Sun Feb 20 13:29:29 2022  | Hide | Hide all
2022-02-20_14-28-02-LowEnergy-EyTotal.png
Attachment 6: 2022-02-20_14-25-45-Tape-server-data-written.png  121 kB  Uploaded Sun Feb 20 13:29:39 2022  | Hide | Hide all
2022-02-20_14-25-45-Tape-server-data-written.png
ELOG V3.1.4-unknown